# Information Discovery in e-Commerce

### v1

@@ 1.1 Motivation
Over the past 20 years, we have seen the explosive growth of e-commerce portals, such as Alibaba, Amazon, eBay, and JD.com. These developments have reshaped people’s’ shopping habits. Accordingly, an increasing number of customers now prefer to spend more time shopping online, generating billions of user requests per day. As part of the process of serving customer requests, large volumes of multi-modal data, including user search logs, clicks, orders, reviews, images, and chat logs, etc., are being generated. From an information retrieval point of view, discovering and employing pertinent information from the sheer volume of e-commerce data so as to enhance the performance of e-commerce services presents interesting challenges, both for academic and industrial researchers. In this survey we describe those challenges and the solutions that the community has so far proposed.
@@
  • e-commerce presentations and users;
@@
E-commerce information discovery problems are wide in scope as the underlying discovery tasks concern a broad range of interaction modalities. There is a growing body of established methods in the area, aimed at developing algorithms for analyzing user behavior, for product search in e-commerce, for recommender systems, and for question answering and dialogue systems. These areas, and the methods developed, form the core around which most ongoing research efforts concerning information discovery for e-commerce are organized. The time is right to organize this material and to present it to a broad audience of interested information retrieval researchers, whether junior or senior, whether academic or industrial [tsagkias-2020-challenges].
@@ 1.3.1 Topics covered
In this survey, we break the e-commerce information discovery problem down into five research directions: (i) e-commerce information presentation and users, (ii) user behavior and profiling in e-commerce, (iii) search in e-commerce, (iv) recommendation in e-commerce, and (v) question answering and dialogue systems in e-commerce. Below, we briefly describe each of these five directions, respectively.
@@
The third direction of this survey concerns search in e-commerce, which examines approaches for product search scenarios on e-commerce portals. Just like, e.g., traditional web search, the target of this task is to satisfy users’ needs. However, product search in e-commerce sites should be realized with different types of features than, e.g., web search, with the availability of a large number of product, query attributes, and engagement features. Moreover, calculating of relevance in product search confronts more challenges about gaps between users and products. Additionally, the target corpora can be structured, semi-structured, or unstructured, or a mixture of these; semantic search against such diverse sources raises interesting research challenges.
The fourth direction concerns recommendations in e-commerce. In contrast to traditional research on recommender systems that focuses on rating prediction, e-commerce recommender systems aim to tackle three challenges: the huge volume of products, sparsity, and data richness. Due to the existence of a very large number of candidate items in e-commerce portals, of which only a small fraction will attract a user’s attention, e-commerce recommendation methods usually follow a two-stage recommendation framework that contains two phases: (i) candidate retrieval, and (ii) candidate ranking. The first phase of candidate retrieval goes through the whole product catalog, and selects a small set of products that might match the information need. The second phase of candidate ranking ranks the candidates to present the final top-$K$ products to the user. Given structured user behavior logs and semi-structured data about product features, e-commerce knowledge bases can be created to assist the candidate generation step. And the candidate ranking procedure ranks the retrieved candidate items for a better conversion rate or click-through rate, based on various machine learning models.
@@ 1.3.2 Topics not covered
E-commerce impacts large parts of our economy and society, including markets and retailers, supply chain management, and employment. With the development of data science, business intelligence studies on e-commerce marketing, e.g., sales volume forecast and time series analysis, are receiving an increasing amount of attention. All of these areas are important, scientifically challenging, and deserving of attention from the information retrieval community. However, our focus will be limited to information discovery within the context of e-commerce. Specifically, we will not address topics such as computational advertising approaches that are irrelevant to search and recommendation, marketing strategies, forecasting, or information management in e-commerce.
@@ 1.3.3 Structure of the survey
The rest of this survey is organized as follows. Chapter 2 provides key definitions and background related to e-commerce information discovery, drawing from user modeling, search, recommender systems, question-answering, and dialogue systems. Chapter 3 describes preliminaries of e-commerce presentations as well as e-commerce users, including user behavior characteristics, and relevant language technologies and their use in e-commerce applications; Chapter 4 details user behavior modeling and user profiling approaches in e-commerce, including click behavior tracking, post-click tracking, purchase behavior modeling, and user profiling in e-commerce. Chapter 5 describes recent approaches proposed for e-commerce search, which we organize along two lines: research about the matching problem in e-commerce search, and about ranking strategies for e-commerce search. Chapter 6 presents algorithms and solutions for recommender systems in e-commerce. After introducing the two-stage recommendation framework in e-commerce portals, we organize the e-commerce recommendation studies into two groups: candidate retrieval models and candidate ranking models. We survey e-commerce question answering and dialogue systems in Chapter 7, where we introduce recent studies on e-commerce question answering and dialogue systems, respectively. In Chapter 8 we conclude this survey and identify emerging research directions and issues for future work.
@@ 1.4 Our readers
We expect this survey to be useful for both academic and industrial researchers who either want to develop e-commerce information discovery methods, use them in their own research, or apply the methods described in the survey to improve product performance in e-commerce services. The intention is to help our audience acquire domain knowledge and to promote information discovery research activities in e-commerce.
To be able to benefit from this survey, we expect the reader to have a background in information retrieval, natural language processing, or machine learning. We recommend that readers read the material that we offer from start to finish, in the order that we offer it. However, readers who have a specific interest in search, or in recommender systems, or in conversational technology in e-commerce should read Chapters 3 and 4 first before skipping ahead to Chapter 5, 6, or 7, respectively.
@@
The chapter presents definitions and background applied to e-commerce information discovery studies from the perspectives of research communities on user modeling, search, recommender systems, quesiton-answering, and dialogue systems. Specifically, we first introduce background-relevant concepts about user modeling, information retrieval, recommender systems, and conversational AI. Thereafter, we introduce definitions and notations associated with e-commerce information discovery. Next, we explore fundamental concepts in e-commerce information discovery, including e-commerce information presentation, e-commerce search, e-commerce recommendation, and e-commerce conversational AI systems. The glossary and notations attached to this concept are introduced in the last part of this chapter.
@@ 2.1 Background
E-commerce has revolutionized how consumers interact with products and services, fundamentally altering the landscape of information discovery. There are plenty of relevant research perspectives on information discovery in e-commerce. Unlike traditional retailing settings, where physical exploration and interaction drive decision-making, e-commerce relies on digital mechanisms to guide users through vast and often overwhelming amounts of information. As a result, the effectiveness of e-commerce platforms hinges on their ability to deliver personalized, relevant, and timely information to users.
Various research disciplines—such as user modeling, information retrieval, recommender systems, question-answering, and conversational AI—contribute significantly to enhancing the e-commerce experience. Each of these fields offers unique insights and methodologies for addressing key challenges in e-commerce, such as understanding user intent, predicting user preferences, addressing user concerns, satisfying user needs, and facilitating seamless product discovery. Targeting on guiding the following chapters, we list the fundamental concepts and research perspectives behind these areas, laying the groundwork for understanding how e-commerce platforms enable efficient and effective information discovery for users.
Information discovery in e-commerce. In the context of e-commerce, information discovery refers to the process by which users engage with relevant products or services based on their specific needs, regardless of the format or presentation of that information. This process encompasses a variety of functions, including search, recommendation, and personalized content delivery and presentation. At its core, information discovery in e-commerce involves not only retrieving relevant products but also understanding user intent and preferences to provide the most suitable results. It relies on sophisticated algorithms to identify, search, recommend, and display information that aligns with user requirements. Whether through search queries, personalized recommendations, or curated content, information discovery systems enable users to efficiently navigate large product catalogs and find what they need in a seamless, engaging manner.
@@ 2.2 User modeling
User modeling refers to the process of creating a representation of a user’s characteristics, behaviors, preferences, and goals in order to personalize user-system interactions or appropriate content for that user. It is widely used in various fields like information retrieval, recommender systems, conversational AI, etc. User modeling is a critical component of personalizing e-commerce experiences. It involves the construction of user profiles based on behavioral data, preferences, demographics, and interactions within the system. These profiles help systems adapt content, recommendations, and interactions to suit individual needs. Meanwhile, e-commerce platforms present information in various formats, such as lists, grids, or interactive elements, which can significantly affect user engagement and conversion rates. Understanding user behavior and preferences by optimizing these presentations is also important for enhancing user experience and satisfaction.
In e-commerce, user modeling can leverage implicit feedback (e.g., clicks, purchases, carting, and user engagement) and explicit feedback (e.g., ratings and reviews) to predict the user future behaviors, preferences, and profiling. Techniques like collaborative filtering, content-based filtering, and hybrid models are frequently employed in user modeling during early studies on this topic. In recent years, deep neural networks and pre-trained language models have been successfully applied to user modeling in e-commerce portals.From a broader research perspective, user modeling spans across domains such as cognitive science, machine learning, and human-computer interaction, contributing to the development of systems that continuously refine the understanding of users as they interact with the platform.
@@ 2.3 Information retrieval in e-commerce
E-commerce search. E-commerce search involves techniques and algorithms used to allow users to efficiently find products or services within an online store. This includes the use of keywords, filters, and advanced semantic search technologies. The e-commerce search differs from traditional information retrieval because it often focuses on product features, pricing, availability, and user preferences. Research in this area spans areas like query understanding, ranking algorithms, and the integration of multimodal data (e.g., images and reviews).
@@
E-commerce recommendations. Moreover, IR in e-commerce is increasingly intertwined with recommender systems. Recommender systems are a cornerstone of e-commerce platforms, helping users discover products they might not have explicitly searched for but are likely to find appealing. These systems predict user preferences using collaborative filtering, content-based filtering, or hybrid methods that combine both approaches. While search engines retrieve items explicitly requested by the user (i.e., with queries), recommendations anticipate user needs by suggesting products the user may not have thought to search for. These systems analyze user data, such as browsing history, purchase patterns, and interactions, to suggest relevant items. They typically use a combination of techniques, including collaborative filtering, which recommends products based on the behaviors of similar users, and content-based filtering, which suggests items with attributes similar to those the user has shown interest in. Many modern systems employ hybrid models that integrate both methods, sometimes enhanced with advanced techniques like deep learning, to improve accuracy and diversity. By delivering personalized recommendations, these systems not only enhance user experience but also drive business goals by increasing engagement, conversion rates, and customer satisfaction, all while introducing users to new products that may surprise or delight them.
@@
Recommendation systems in e-commerce analyze user data and behavioral patterns to suggest products or services that users are likely to be interested in. These systems utilize various algorithms, including collaborative filtering, content-based filtering, and hybrid approaches. In the context of e-commerce, recommender systems must balance relevance, diversity, novelty, and serendipity to enhance user engagement and satisfaction. By dividing into candidate retrieval and reranking stages, these systems often operate in two modes: personalized recommendations (based on individual profiles and history) and non-personalized recommendations (based on overall product popularity or trends). Research in recommender systems for e-commerce involves improving recommendation algorithms, addressing challenges like cold-start users, and optimizing recommendations for business goals such as conversion rates and customer retention.
@@
In summary, information retrieval is foundational to the search and recommendation functions within e-commerce platforms. This convergence between search and recommendation highlights the importance of IR techniques that can balance precision (retrieving highly relevant products) with recall (offering a broader set of options that might interest the user). Hybrid models that combine collaborative filtering, content-based filtering, and neural IR approaches are commonly employed to address this dual need. The integration of advanced IR techniques, such as search and recommendation models, allows e-commerce platforms to deliver highly personalized, efficient, and contextually relevant user experiences, ensuring that users find the products they want—and even those they didn’t know they wanted.
@@ 2.4 Conversational AI
E-commerce question-answering. Question-answering (QA) systems are designed to deliver direct and precise responses to user questions, improving both user satisfaction and decision-making efficiency. Recently, more and more e-commerce platforms start to provide question-answering services. E-commerce QA systems help to enhance user experiences by enabling customers to obtain relevant, concise, and accurate answers to their product-related queries. These systems typically understand user intent and either retrieve answers from existing product descriptions, reviews, and FAQs or generate responses dynamically using advanced models. E-commerce question-answering refers to the process in which users ask product-related questions on an e-commerce platform, and the system provides answers either from existing knowledge bases, reviews, or user-generated content. E-commerce QA systems leverage both retrieval-based and generative models to match or generate appropriate answers. This helps users make informed decisions based on product descriptions, user reviews, and frequently asked questions (FAQs). The goal is to reduce information overload and improve the user experience by providing relevant and concise answers.
@@
E-commerce QA systems must handle a wide variety of queries ranging from simple fact-based questions (e.g., "What is the price of this product?") to more complex inquiries about product specifications, reviews, or usage (e.g., "Is this laptop suitable for gaming?"). To address this diversity, QA systems often incorporate a mix of retrieval-based approaches, which search for relevant information in structured data or knowledge bases, and generative approaches, which generate answers when information is sparse or not directly available. Additionally, many e-commerce platforms enable community-based QA, where previous buyers or users of a product can contribute answers, further enriching the system’s knowledge base. The integration of QA systems into e-commerce portals helps reduce the friction often associated with product discovery and decision-making. By offering immediate answers to user queries, these systems improve the overall shopping experience, increase user engagement, and can positively impact conversion rates. E-commerce QA is a rapidly evolving field with ongoing research aimed at improving the accuracy, efficiency, and personalization of responses.
@@
E-commerce automatic dialogue systems. Automatic dialogue systems aim to engage in natural, human-like conversations with users and are widely used in applications such as customer support, virtual assistants, and e-commerce. They provide personalized, efficient, and engaging interactions, enhancing the overall user experience. In e-commerce, automatic dialogue systems, often in the form of chatbots or voice assistants, enable natural language interactions between users and platforms. They can support multi-turn interactions, where users ask follow-up questions, refine their preferences, or seek assistance, creating a more engaging and personalized shopping experience. These systems assist users in discovering information, finding products, completing purchases, and sharing their opinions, all through conversational interfaces. Systematically, automatic dialogue systems in e-commerce refers to the use of chatbots and virtual assistants that can simulate a human conversation to assist users in finding products, answering inquiries, and facilitating transactions. These systems leverage natural language processing and machine learning to provide timely and relevant assistance.
@@
Towards both single-turn and multi-turn interactions between users and systems, research in conversational AI for e-commerce focuses on improving dialogue understanding, response generation, context retention across sessions, and user satisfaction. Additionally, conversational AI systems must adapt to diverse user needs and accommodate various languages and cultural contexts, making this an evolving area of study.
@@
We divide this chapter into two sections: e-commerce presentations (Section 3.1) and e-commerce users (Section 3.2). In Section 3.1, we first introduce basic concepts of e-commerce interfaces; then we detail studies that analyze different aspects of e-commerce presentations, i.e., title analysis, item information analysis, and review analysis. In Section 3.2, we list characteristics of e-commerce users, and examine user behavior on e-commerce portals, i.e., macro behavior, micro behavior and cross-platform behavior.

### v3

@@ 1.1 Motivation
Over the past 20 years, we have seen an explosive growth of e-commerce portals, such as Alibaba, Amazon, eBay, and JD.com. These developments have reshaped people’s shopping habits. An increasing number of customers now prefer to spend more time shopping online, generating billions of user requests per day. As part of the process of serving customer requests, large volumes of multi-modal data, including user search logs, clicks, orders, reviews, images, and chat logs, etc., are being generated. From an information retrieval point of view, discovering and employing pertinent information from the sheer volume of e-commerce data so as to enhance the performance of e-commerce services presents interesting challenges, both for academic and industrial researchers. In this survey we describe those challenges and the solutions that the community has so far proposed.
@@
  • e-commerce presentation and users;
@@
E-commerce information discovery problems are wide in scope as the underlying discovery tasks concern a broad range of interaction modalities. There is a growing body of established methods in the e-commerce, aimed at developing algorithms for analyzing user behavior, for product search, for recommender systems, and for question answering and dialogue systems. These areas, and the methods developed, form the core around which most ongoing research efforts concerning information discovery for e-commerce are organized. The time is right to organize this material and to present it to a broad audience of interested information retrieval researchers, whether junior or senior, whether academic or industrial [tsagkias-2020-challenges].
@@ 1.3.1 Topics covered
We break the e-commerce information discovery problem down into five research directions: (i) e-commerce information presentation and users, (ii) user behavior and profiling in e-commerce, (iii) search in e-commerce, (iv) recommendation in e-commerce, and (v) question answering and dialogue systems in e-commerce. Below, we briefly describe each of these five directions.
@@
The third direction of this survey concerns search in e-commerce, which examines approaches for product search scenarios on e-commerce portals. Just like, e.g., traditional web search, the target of this task is to satisfy users’ needs. However, product search in e-commerce sites should be realized with different types of features than, e.g., web search, with the availability of a large number of product, query attributes, and engagement features. Moreover, calculating relevance in product search faces challenges regarding gaps between users and products. The target corpora can be structured, semi-structured, or unstructured, or a mixture of these; semantic search against such diverse sources raises interesting research challenges.
The fourth direction concerns recommendations in e-commerce. In contrast to traditional research on recommender systems that focuses on rating prediction, e-commerce recommender systems aim to tackle three challenges: the huge volume of products, sparsity, and data richness. Due to the existence of a very large number of candidate items in e-commerce portals, of which only a small fraction will attract a user’s attention, e-commerce recommendation methods usually follow a two-stage recommendation framework with (i) candidate retrieval, and (ii) candidate ranking. The first phase of candidate retrieval goes through the whole product catalog, and selects a small set of products that might match the information need. The second phase of candidate ranking ranks the candidates to present the final top-$K$ products to the user. Given structured user behavior logs and semi-structured data about product features, e-commerce knowledge bases can be created to assist the candidate generation step. And the candidate ranking procedure ranks the retrieved candidate items for a better conversion rate or click-through rate, based on various machine learning models.
@@ 1.3.2 Topics not covered
E-commerce impacts large parts of our economy and society, including markets and retailers, supply chain management, and employment. With the development of data science, business intelligence studies on e-commerce marketing, e.g., sales volume forecasting and time series analysis, are receiving an increasing amount of attention. All of these areas are important, scientifically challenging, and deserving of attention from the information retrieval community. However, our focus will be limited to information discovery within the context of e-commerce. Specifically, we will not address topics such as computational advertising approaches that are irrelevant to search and recommendation, marketing strategies, forecasting, or information management in e-commerce.
@@ 1.3.3 Structure of the survey
The remainder of this survey is organized as follows. Section 2 provides key definitions and background related to e-commerce information discovery, drawing from user modeling, search, recommender systems, question-answering, and dialogue systems. Section 3 describes preliminaries of e-commerce presentations as well as e-commerce users, including user behavior characteristics, and relevant language technologies and their use in e-commerce applications. Section 4 details user behavior modeling and user profiling approaches in e-commerce, including click behavior tracking, post-click tracking, purchase behavior modeling, and user profiling in e-commerce. Section 5 describes recent approaches proposed for e-commerce search, which we organize along two lines: research about the matching problem in e-commerce search, and about ranking strategies for e-commerce search. Section 6 presents algorithms and solutions for recommender systems in e-commerce. After introducing the two-stage recommendation framework in e-commerce portals, we organize the e-commerce recommendation studies into two groups: candidate retrieval models and candidate ranking models. We survey e-commerce question answering and dialogue systems in Section 7, where we introduce recent studies on e-commerce question answering and dialogue systems, respectively. In Section 8 we conclude this survey and identify emerging research directions and issues for future work.
@@ 1.4 Our readers
We expect this survey to be useful to both academic and industrial researchers who either want to develop e-commerce information discovery methods, use them in their own research, or apply the methods described in the survey to improve product performance in e-commerce services. The intention is to help our audience acquire domain knowledge and to promote information discovery research activities in e-commerce.
To be able to benefit from this survey, we expect the reader to have a background in information retrieval, natural language processing, or machine learning. We recommend that readers read the material that we offer from start to finish, in the order that we offer it. However, readers who have a specific interest in search, or in recommender systems, or in conversational technology in e-commerce should read Section 3 and 4 first before skipping ahead to Section 5, 6, or 7, respectively.
@@
The section presents definitions and background applied to e-commerce information discovery studies from the perspectives of research communities on user modeling, search, recommender systems, quesiton-answering, and dialogue systems. We first introduce relevant concepts about user modeling, information retrieval, recommender systems, and conversational AI. We then introduce definitions and notations associated with e-commerce information discovery. Next, we explore fundamental concepts in e-commerce information discovery, including e-commerce information presentation, e-commerce search, e-commerce recommendation, and e-commerce conversational AI systems. The glossary and notations attached to these concepts are introduced in the last part of this section.
@@ 2.1 Background
E-commerce has revolutionized how consumers interact with products and services, fundamentally altering the landscape of information discovery. There are plenty of relevant research perspectives on information discovery in e-commerce. Unlike traditional retail settings, where physical exploration and interaction drive decision-making, e-commerce relies on digital mechanisms to guide users through vast and often overwhelming amounts of information. As a result, the effectiveness of e-commerce platforms hinges on their ability to deliver personalized, relevant, and timely information to users.
Various research disciplines – such as user modeling, information retrieval, recommender systems, question-answering, and conversational AI – contribute significantly to enhancing the e-commerce experience. Each of these fields offers unique insights and methodologies for addressing key challenges in e-commerce, such as understanding user intent, predicting user preferences, addressing user concerns, satisfying user needs, and facilitating seamless product discovery. We list the fundamental concepts and research perspectives behind these areas, laying the groundwork for understanding how e-commerce platforms enable efficient and effective information discovery for users.
Information discovery in e-commerce. In the context of e-commerce, information discovery refers to the process by which users engage with relevant products or services based on their specific needs, regardless of the format or presentation of that information. This process encompasses a variety of functions, including search, recommendation, and personalized content delivery and presentation. At its core, information discovery in e-commerce involves not only retrieving relevant products but also understanding user intent and preferences to provide the most suitable results. It relies on algorithmic solutions to identify, search, recommend, and display information that aligns with user requirements. Whether through search queries, personalized recommendations, or curated content, information discovery systems enable users to efficiently navigate large product catalogs and find what they need in a seamless and engaging manner.
@@ 2.2 User modeling
User modeling refers to the process of creating a representation of a user’s characteristics, behaviors, preferences, and goals in order to personalize user-system interactions or appropriate content for that user. It is widely used in fields like information retrieval, recommender systems, and conversational AI. User modeling is a critical component of personalizing e-commerce experiences. It involves the construction of user profiles based on behavioral data, preferences, demographics, and interactions within the system. These profiles help systems adapt content, recommendations, and interactions to suit individual needs. Meanwhile, e-commerce platforms present information in various formats, such as lists, grids, or interactive elements, which can significantly affect user engagement and conversion rates. Understanding user behavior and preferences by optimizing these presentations is important for enhancing the user experience and user satisfaction.
In e-commerce, user modeling can use implicit feedback (e.g., clicks, purchases, carting, and user engagement) and explicit feedback (e.g., ratings and reviews) to predict a user’s future behavior, preferences, and profiling. Techniques like collaborative filtering, content-based filtering, and hybrid models are frequently employed in user modeling during early studies on this topic. In recent years, deep neural networks and pre-trained language models have been successfully applied to user modeling in e-commerce portals. User modeling spans across domains such as cognitive science, machine learning, and human-computer interaction, contributing to the development of systems that continuously refine the understanding of users as they interact with the platform.
@@ 2.3 Information retrieval in e-commerce
E-commerce search. E-commerce search involves techniques and algorithms used to allow users to efficiently find products or services within an online store. This includes the use of keywords, filters, and advanced semantic search technologies. E-commerce search differs from traditional information retrieval because it often focuses on product features, pricing, availability, and user preferences. Research in this area spans areas like query understanding, ranking algorithms, and the integration of multimodal data (e.g., images and reviews).
@@
E-commerce recommendations. IR in e-commerce is increasingly intertwined with recommender systems. Recommender systems are a cornerstone of e-commerce platforms, helping users discover products they might not have explicitly searched for but are likely to find appealing. These systems predict user preferences using collaborative filtering, content-based filtering, or hybrid methods that combine both approaches. While search engines retrieve items explicitly requested by the user (i.e., through queries), recommendations anticipate user needs by suggesting products the user may not have thought to search for. These systems analyze user data, such as browsing history, purchase patterns, and interactions, to suggest relevant items. They typically use a combination of techniques, including collaborative filtering, which recommends products based on the behaviors of similar users, and content-based filtering, which suggests items with attributes similar to those the user has shown interest in. Many modern systems employ hybrid models that integrate both types of method, sometimes enhanced with techniques like deep learning, to improve accuracy and diversity. By delivering personalized recommendations, these systems not only enhance the user experience but also drive business goals by increasing engagement, conversion rates, and customer satisfaction, all while introducing users to new products that may surprise or delight them.
@@
Recommendation systems in e-commerce analyze user data and behavioral patterns to suggest products or services that users are likely to be interested in. These systems use various algorithms, including collaborative filtering, content-based filtering, and hybrid approaches. In the context of e-commerce, recommender systems must balance relevance, diversity, novelty, and serendipity to enhance user engagement and satisfaction. By using a division into candidate retrieval and reranking stages, these systems often operate in two modes: personalized recommendations (based on individual profiles and history) and non-personalized recommendations (based on overall product popularity or trends). Research in recommender systems for e-commerce involves improving recommendation algorithms, addressing challenges like cold-start users, and optimizing recommendations for business goals such as conversion rates and customer retention.
@@
In summary, information retrieval is foundational to the search and recommendation functions within e-commerce platforms. The convergence between search and recommendation highlights the importance of IR techniques that can balance precision (retrieving highly relevant products) with recall (offering a broader set of options that might interest the user). Hybrid models that combine collaborative filtering, content-based filtering, and neural IR approaches are commonly employed to address this dual need. The integration of advanced IR techniques, such as search and recommendation models, allows e-commerce platforms to deliver highly personalized, efficient, and contextually relevant user experiences, ensuring that users find the products they want – and even those they did not know they wanted.
@@ 2.4 Conversational AI
E-commerce question-answering. Question-answering (QA) systems are designed to deliver direct and precise responses to user questions, improving both user satisfaction and decision-making efficiency. Recently, e-commerce platforms have started to provide question-answering services. E-commerce QA systems help to enhance user experiences by enabling customers to obtain relevant, concise, and accurate answers to their product-related queries. These systems typically understand user intent and either retrieve answers from product descriptions, reviews, and FAQs or generate responses dynamically using advanced models. E-commerce question-answering refers to the process in which users ask product-related questions on an e-commerce platform, and the system provides answers either from knowledge bases, reviews, or user-generated content. E-commerce QA systems use both retrieval-based and generative models to match or generate appropriate answers. This helps users make informed decisions based on product descriptions, user reviews, and frequently asked questions (FAQs). The goal is to reduce information overload and improve the user experience by providing relevant and concise answers.
@@
E-commerce QA systems must handle a wide variety of queries ranging from simple fact-based questions (e.g., “What is the price of this product?”) to more complex inquiries about product specifications, reviews, or usage (e.g., “Is this laptop suitable for gaming?”). To address this diversity, QA systems often incorporate a mix of retrieval-based approaches, which search for relevant information in structured data or knowledge bases, and generative approaches, which generate answers when information is sparse or not directly available. Additionally, many e-commerce platforms enable community-based QA, where previous buyers or users of a product can contribute answers, further enriching the system’s knowledge base. The integration of QA systems into e-commerce portals helps reduce the friction often associated with product discovery and decision-making. By offering immediate answers to user queries, these systems improve the overall shopping experience, increase user engagement, and can positively impact conversion rates. E-commerce QA is a rapidly evolving field with ongoing research aimed at improving the accuracy, efficiency, and personalization of responses.
@@
E-commerce automatic dialogue systems. Automatic dialogue systems aim to engage in natural, human-like conversations with users and are widely used in applications such as customer support, virtual assistants, and e-commerce. They provide personalized, efficient, and engaging interactions, enhancing the overall user experience. In e-commerce, automatic dialogue systems, often in the form of chatbots or voice assistants, enable natural language interactions between users and platforms. They can support multi-turn interactions, where users ask follow-up questions, refine their preferences, or seek assistance, creating a more engaging and personalized shopping experience. These systems assist users in discovering information, finding products, completing purchases, and sharing their opinions, all through conversational interfaces. Systematically, automatic dialogue systems in e-commerce refers to the use of chatbots and virtual assistants that can simulate a human conversation to assist users in finding products, answering inquiries, and facilitating transactions. These systems use natural language processing and machine learning to provide timely and relevant assistance.
@@
Research in conversational AI for e-commerce focuses on improving dialogue understanding, response generation, context retention across sessions, and user satisfaction. Additionally, conversational AI systems must adapt to diverse user needs and accommodate various languages and cultural contexts, making this an evolving area of study.
@@
We divide this section into two parts: e-commerce presentations (Section 3.1) and e-commerce users (Section 3.2). In Section 3.1, we first introduce basic concepts of e-commerce interfaces; then we detail studies that analyze different aspects of e-commerce presentations, i.e., title analysis, item information analysis, and review analysis. In Section 3.2, we list characteristics of e-commerce users, and examine user behavior on e-commerce portals, i.e., macro behavior, micro behavior and cross-platform behavior.